### v1

# Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registrations

Abstract
In this paper, we discuss Mahalanobis k-NN: a statistical lens designed to address the challenges of feature matching in learning-based point cloud registration when confronted with an arbitrary density of point clouds, either in the source or target point cloud. We tackle this by adopting Mahalanobis k-NN’s inherent property to capture the distribution of the local neighborhood and surficial geometry. Our method can be seamlessly integrated into any local-graph-based point cloud analysis method. In this paper, we focus on two distinct methodologies: Deep Closest Point (DCP) and Deep Universal Manifold Embedding (DeepUME). Our extensive benchmarking on the ModelNet40 and Faust datasets highlights the efficacy of the proposed method in point cloud registration tasks. Moreover, we establish for the first time that the features acquired through point cloud registration inherently can possess discriminative capabilities. This is evident by a substantial improvement of about 20% in the average accuracy observed in the point cloud few-shot classification task benchmarked on ModelNet40 and ScanObjectNN. The code is publicly available at https://github.com/TejasAnvekar/Mahalanobis-k-NN
Keywords:
Point Cloud Registration Metric Distance 3D Vision.
Figure 1: The visual supremacy of our proposed methodologies, MDCP-v1 and MDeepUME, becomes apparent in a point cloud registration task involving a target point cloud with only half the points compared to the source point cloud. In the visualization, highlighted boxes in red illustrate the limited performance of DCP-v1 [DCP] and DeepUME [DeepUME], while the highlighted boxes in green demonstrate the resilience of the proposed Mahalanobis versions of DCP-v1 [DCP] and DeepUME [DeepUME].
1 Introduction
With the rapid progress of high-precision sensors like LiDAR [lidar] and Kinect [kinect], point clouds have become the prevalent data format for 3D representation. Due to sensor limitations in capturing only partial views, registration algorithms are crucial for amalgamating scans into comprehensive 3D scenes. Point cloud registration involves determining transformation matrices between point cloud pairs, facilitating the fusion of partial scans to produce coherent 3D representations.
Registration is a fundamental task spanning various computational fields such as medical imaging [medicalImaging], robotics [robotics], autonomous driving [autoDriving], and computational chemistry [ComChe]. Estimating transformation matrices serves applications like 3D reconstruction [reconstruction] and precise 3D localization [Localization]. It underpins the development of intricate 3D maps for autonomous driving [autoDriving], environment reconstruction in robotics [robotics], and improved safety in mining. Additionally, registration’s ability to facilitate high-precision localization proves invaluable for entities like driverless cars [autoDriving], ensuring precise positioning and interaction within the 3D environment.
The registration of point clouds has garnered significant attention through closed-form [ICP, goICP, FGR, aligningPointCloudView, ume] and learning-based [pointnetLK, DCP, DeepGMR, DeepUME, rgm] methodologies. DCP [DCP] tackles the challenge of feature matching (correspondences) by initially estimating local features via weight-shared DGCNN [dgcnn] (Dynamic Graph Convolutional Neural Network). Subsequent steps involve an attention-based module and differentiable SVD modules for point-to-point registration. Conversely, DeepUME [DeepUME] adopts a distinct approach by projecting points into a space invariant under $\mathcal{SO}(3)$ transformations. This space is leveraged to compute invariant per-point features that address real-world data sampling issues. Employing two weight-shared transformers and DGCNN [dgcnn], correspondences are established and fed into UME [ume] for rigid transformation parameter estimation. A limitation of both DCP [DCP] and DeepUME [DeepUME] is susceptibility to failure in scenarios where either the source or target point cloud exhibits arbitrary (lower) density than the others—an occurrence commonly encountered in 3D scan registration from diverse sensors. This limitation is attributed to the methods’ reliance on estimating per-point correspondences using local features extracted from edge-conv [dgcnn], which operates on graphs constructed using k-nearest neighbors (k-NN).
We introduce Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration to mitigate this issue. The efficacy of Mahalanobis distance as an evaluation metric has already been demonstrated in reference-based point cloud quality assessment [MD1]. Our proposed Mahalanobis k-NN can be used as a plugin for any point-cloud registration method. In this paper, we propose mahalnobis versions of DCP-(v1,v2) [DCP] and DeepUME [DeepUME]. Both methods operate on diverse point-cloud representation spaces. This augmentation facilitates pointcloud registration as shown in Figure 1; across various publicly available 3D datasets, encompassing diverse benchmarking scenarios, including: 1) unseen-category evaluation, 2) robustness to various noise types, and 3) efficiency towards varying point densities. Additionally, we demonstrate the discriminative capabilities of the proposed Mahalanobis lens through point cloud few-shot classification tasks. In this context, models pre-trained for registration tasks leverage DGCNN [dgcnn] features for few-shot evaluation on both ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets. Finally, we summarize our contributions as follows:
  • We propose Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration by incorporating Mahalanobis distance in:
    • Deep Closest Point (DCP) [DCP] that operates on Euclidean coordinate space.
    • Deep Universal Manifold Embedding (DeepUME) [DeepUME] that operates on $\mathcal{SO}(3)$ invarient space.
  • We demonstrate the efficacy of the proposed method on various point cloud registration tasks on a variety of publicly available data sets and compare the results with state-of-the-art techniques.
  • We perform an endurance test to evaluate the robustness and generalization of the proposed method. We achieve state-of-the-art results compared to other point cloud registration methods on all benchmarking strategies and on almost all evaluation metrics.
  • To the best of our knowledge, we are the first to demonstrate discriminative prowess inherited in point cloud registration task by benchmarking point cloud few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets through features extracted by DCP [DCP], we report up to an average of 20% increment when Mahalanobis is incorporated.
  • We will release our code to facilitate reproducibility and future research upon acceptance.
2 Related Works
Optimization Based. Preceding the advent of the deep learning era in 3D point cloud registration, one common strategy involved extracting and matching spatially local features, as demonstrated in studies like [comprehensiveEvalFeatureDescriptor, spinImages, aligningPointCloudView, RepeatabilityRobust3Dmatching, fastRobustDescriptor, rotationalContour]. Many existing methods in this category are adaptations of 2D image processing solutions, such as variants of 3D-SIFT [3Dsift] and the 3D Harris key-point detector [harris3D]. However, key-point matching in 3D presents challenges due to the absence of a regular sampling grid, artifacts, and sampling noise, leading to high outlier rates and localization errors. To achieve global alignment, key-point matching typically utilizes outlier rejection methods like RANSAC [ransac] and is refined using local optimization algorithms [ICP, scanRegistration3Dndt, registrationWithoutICP, iterativePointMatching]. Notably, DGR [DGR] follows a similar paradigm, but it incorporates learnable inlier detection. Researchers have proposed numerous works for handling outliers and noise [trimmedICP], formulating robust minimizers [robustRegistration2D3D], and devising more suitable distance metrics. The widely used Iterative Closest Point algorithm (ICP) [ICP, iterativePointMatching], a popular refinement algorithm, constructs point correspondences based on spatial proximity and employs a transformation estimation step. Over time, various ICP variants [efficientICP, goICP, sparseICP] have been proposed to enhance convergence rate, robustness, and accuracy in the 3D point cloud registration.
Learning Based. Registration approaches leverage learning-based techniques alongside closed-form methods. Pioneered by PointNet [pointnet] and advanced by DGCNN [dgcnn], learning from data for point cloud registration has become powerful (e.g., PointNet++, DCP [DCP], PRNet [prnet]). PointNetLK [pointnetLK] minimizes learned feature distance using a differentiable Lucas-Kanade algorithm, while DCP uses attention and differentiable SVD for point-to-point registration. Recent work [rgm] introduces graph transformation and deep graph matching. Moreover, DeepUME [DeepUME] employ $\mathcal{SO}(3)$ invariant coordinate systems for global registration. DeepGMR [DeepGMR] tackles pose-invariant correspondences between raw point clouds and Gaussian-mixture models but suffers from noise sensitivity. DeepUME [DeepUME] fuses closed-form Universal Manifold Embedding (UME) and deep learning to handle sparse and unevenly sampled point clouds with large transformations. This framework is trained end-to-end and learns joint-resampling strategies and invariant features.
DCP [DCP] and DeepUME [DeepUME] achieve impressive performance but struggle when the source or target point cloud densities are unequal. Both rely on DGCNN [dgcnn] for local feature-based correspondence estimation. However, DGCNN’s Euclidean metric for k-nearest neighbors lacks surface awareness. We propose Mahalanobis versions of DCP and DeepUME, introducing the Mahalanobis distance for surface-aware features and enhanced registration in real-world scenarios. Furthermore, our proposed approach demonstrates remarkable discriminative capabilities, as evident from its performance in point cloud few-shot classification tasks.
3 Problem Setting
We denote $x_{i}^{T}(i\in[1,M])$ and $y_{i}^{T}(i\in[1,N])$ as row vectors from matrices $X\in\mathbb{R}^{M\times 3}$ and $Y\in\mathbb{R}^{N\times 3}$ respectively. The matrices $X$ and $Y$ represent two distinct point clouds, often originating from two scans of the same object (we assume $Y$ is transformed from $X$ by an unknown rigid motion). Here, $x_{i}$ and $y_{j}$ represent the coordinates of the $i$th point in $X$ and the $j$th point in $Y$, for $i\in[1,M]$ and $j\in[1,N]$ respectively. Assuming that there are $S$ pairs of corresponding points between the two point clouds, the goal of registration is to determine the optimal rigid transformation parameters $g$ (comprising a rotation matrix $R\in\mathcal{SO}(3)$ and a translation vector $t\in\mathbb{R}^{3}$) that aligns point cloud $X$ to point cloud $Y$, as illustrated below:
$\operatorname*{arg\,min}_{R\in\mathcal{SO}(3),\,t\in\mathbb{R}^{3}}\left\|d\left(X,g(Y)\right)\right\|_{2}^{2}$ (1)
The term $d\big{(}X,g(Y)\big{)}$ is a measure of the projection error between $X$ and the transformed $Y$, expressed as $d\big{(}X,RY+t\big{)}=\sum_{s=1}^{S}\left\|x_{s}-\big{(}Ry_{s}+t\big{)}\right\|_{2}$, where $s$ iterates over the $S$ pairs of corresponding points. The optimization problem in Equation 1 embodies a well-known “chicken-and-egg” scenario: determining the optimal transformation matrix requires knowledge of true correspondences  [correspondance1, correspondance2]; conversely, accurate correspondences can be established given the optimal transformation matrix. However, the simultaneous resolution of both aspects poses a non-trivial challenge.
4 Method
In this paper, we focus on point cloud registration using DCP [DCP] and DeepUME [DeepUME] for global one-shot rigid transformations (contrasting iterative methods like ICP [ICP]). Both rely on DGCNN [dgcnn] for local features but use static k-nearest neighbor (k-NN) graphs, limiting semantic space creation. Unlike, DGCNN’s dynamic k-NN graphs, these methods restrict feature projection depth. We propose using the Mahalanobis distance [MD] to enhance k-NN, enabling surface-aware neighbor selection based on principal components. This extracts surface and corner features crucial for accurate matching and robust transformation estimation.
4.1 Mahalanobis k-NN
Figure 2: The illustration contrasts Euclidean distance fields (red) capturing spatial neighbors with Mahalanobis distance fields (blue) considering the underlying data distribution. A chair point cloud exemplifies this. The black query point is surrounded by Euclidean neighbors (red points) and Mahalanobis neighbors (blue points). The depiction clearly illustrates the impact of Mahalanobis distance, effectively capturing surface points per the data distribution vital for precise feature matching.
In this section, we introduce Mahalanobis distance as a statistical lens for robust feature matching in DCP [DCP] and DeepUME [DeepUME]. We modify DGCNN to construct a graph $\mathcal{G}^{\dagger}$ using Mahalanobis k-nearest neighbors (k-NN) for enhanced feature selection. DGCNN constructs a graph $\mathcal{G}$, applies nonlinearity for edge values, and performs vertex-wise aggregation ($\max$ or $\sum$) in each layer. Let $x_{i}^{l}$ be the embedding of point $i$ in the $l$-th layer, and $h_{\theta}^{l}$ be a nonlinear function parameterized by a shared multi-layer perceptron (MLP). DGCNN can be formulated as:
$x^{l}_{i}=f\Big{(}\big{\{}h_{\theta}^{l}(x_{i}^{l-1},x_{j}^{l-1})\forall j\in\mathcal{N}_{i}\big{\}}\Big{)}$ (2)
where $\mathcal{N}_{i}$ denotes neighbors of vertex $i$ in $\mathcal{G}$.
We propose a novel augmentation by integrating Mahalanobis k-nearest neighbors (k-NN) in Equation.2 to construct graph $\mathcal{G}^{\dagger}$. This graph prioritizes surface awareness through selection of surficial neighbors $\mathcal{N}_{i}^{\dagger}$ based on Mahalanobis distance as shown in Figure 2.
To compute Mahalanobis distance between points $x_{i}$ and $x_{j}$, we use:
$D_{M}(x_{i},x_{j})=\sqrt{(x_{i}-x_{j})^{T}\cdot C^{-1}\cdot(x_{i}-x_{j})}$ (3)
where: $D_{M}(x_{i},x_{j})$ is Mahalanobis distance between $x_{i}$ and $x_{j}$, and $C$ is covariance matrix of data points. One can argue, geodesic graph-based methods capture similar or sometimes better surficial information. To counter this, we implement vectorized Floyd-Warshall (refer supplementary for implementation detail on vectorized Floyd-Warshall) on the Euclidean-knn graph and observe that the performance is slower and also not as effective as the proposed Mahalanobis variants of DCP and DeepUME.
The impact of Mahalanobis distance is visually depicted in Figure 2. This surficial awareness proves crucial in feature matching. Incorporating Mahalanobis distance into DCP  [DCP] and DeepUME  [DeepUME] yields superior performance in point cloud registration. Discriminative power is affirmed in point-cloud few-shot classification, with pre-trained DGCNN models from vanilla DCP and Mahalanobis DCP showcasing remarkable classification capabilities.
5 Experiments
In this section, we showcase the superior performance of our proposed method in the domain of point cloud registration when compared against existing state-of-the-art techniques such as ICP [ICP], GO-ICP [goICP], FGR [FGR], PointNetLK [pointnetLK], DCP [DCP], and DeepUME [DeepUME] on publicly available datasets. Our evaluation encompasses a spectrum of benchmarking strategies, encompassing scenarios where one point cloud is notably sparser than the other. For a fair comparison, we implemented a vectorized Floyd-Warshall algorithm in PyTorch to compute geodesic distances and compare it to our proposed method for surface feature extraction. We further validate the Mahalanobis metric’s discriminative power in few-shot learning tasks and showcase robustness through endurance tests. Our model was trained using an Nvidia RTX 3090 GPU and PyTorch 1.11.
5.1 Point Cloud Registration
We build upon the model architecture, training strategies, evaluation metrics, and hyperparameters established in DCP [DCP]  [^1] https://github.com/WangYueFt/dcp. We introduce the Mahalanobis version of DCP (MDCP), detailed in Section 4.1. MDCP comes in two variants: v1 without a transformer and v2 with a transformer. We further extend our approach to DeepUME [DeepUME] by proposing the Mahalanobis variant, MDeepUME. For a fair comparison, we benchmark geodesic versions of these models using vectorized Floyd-Warshall on Euclidean k-NN graphs.
Dataset. To facilitate benchmarking, we employ publicly available datasets: ModelNet40 [modelnet], FAUST [FAUST], and Stanford3D (S-3D) [^2] http:// graphics.stanford.edu/data/3Dscanrep/ . Notably, the latter two datasets are exclusively employed for testing purposes. To ensure methodological consistency, we adhere to the official training/testing splits and settings as stipulated in the original works of DCP [DCP] and DeepUME [DeepUME]. Among these datasets, our approach is rigorously evaluated.
We evaluate using root mean squared error (RMSE) between ground truth and predicted transformations (lower is better). Angular errors are reported in degrees. These metrics comprehensively assess our approach’s performance.
5.2 Comparison with State-of-the-art Methods
Table 1: Comparison on ModelNet40 [modelnet] for unseen data, unseen category, and Gaussian noise. MDCP-(v1,v2) and MDeepUME outperform baselines (bold best, underline second best). Lower is better. $\dagger$ denotes reproduced results, -Geo refers to Floyd-Warshall geodesic-graph version.
Unseen data Unseen Category Unseen data + Noise
RMSE(R) RMSE(t) RMSE(R) RMSE(t) RMSE(R) RMSE(t)
ICP [ICP] 29.9148 0.2909 29.8764 0.2933 29.7080 0.2906
GO-ICP [goICP] 11.8523 0.0257 13.8657 0.0226 11.4535 0.0231
FGR [FGR] 9.3628 0.0139 9.8490 0.0135 24.6515 0.1090
PointNetLK [pointnetLK] 15.0954 0.0221 17.5021 0.0280 16.0049 0.0216
[…]
MDCP-v1 (Ours) 1.6614 0.0004 1.9679 0.0004 1.5889 0.0004
[…]
MDCP-v2 (Ours) 0.9468 0.0004 1.3849 0.0049 0.8481 0.0043
DeepUME [DeepUME] 0.0062 0 0.0183 0.00019 2.5263 0.0006
[…]
MDeepUME 0.0023 0 0.0098 0.00009 1.9946 0.0006
ModelNet40 Results. Table 1 compare our methods (MDCP, MDeepUME) with baselines and their geodesic versions. We outperform all methods across various settings: Unseen data: Train-test split from ModelNet40 [modelnet]; Unseen category: First 20 classes for training, remaining for testing; Gaussian Noise: $\mathcal{N}(0,0.01)$ clipped to [-0.05, 0.05]. The point density is 1024 in all settings. Baseline evaluations are from [DCP] (highly comparable results achieved with minimal standard deviation during our reproduction attempts). For each case, a rigid transformation is applied along each axis (rotation: $[0^{\circ},45^{\circ}]$, translation: [-0.5, 0.5]). The network receives both the original and transformed point cloud as input, evaluated against ground truth using MDCP and MDeepUME.
Figure 3: Robustness evaluation on unseen data (FAUST [FAUST] hand), low-density source (airplane), and low-density target (chair) point clouds (trained on ModelNet40 [modelnet]). Red boxes highlights vulnerabilities in DCP-v1 [DCP], while green boxes highlights MDCP-v1’s efficacy.
Table 2: Evaluation of few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] (OBJ ONLY, OBJ+BG, PB75 splits). Our DCGNN in MDCP-(v1,v2) outperforms baselines across all settings (bold denotes best accuracy). We report mean accuracy (%) and standard deviation from 50 runs.
ModelNet40 [modelnet]
5 way 10 way
Method 10 shot 20 shot 10 shot 20 shot
DCP-v1 [DCP] 67.6 $\pm$6.91 69 $\pm$7.87 50.5 $\pm$7.74 57.1 $\pm$5.01
MDCP-v1 (Ours) 75.2 $\pm$7.3 80 $\pm$7.74 71.0 $\pm$4.16 70.9 $\pm$5.72
Increment in % 11.24 15.94 40.59 24.17
DCP-v2 [DCP] 76.8 $\pm$4.01 79.8 $\pm$6.17 60.8 $\pm$4.93 69.15 $\pm$5.7
MDCP-v2 (Ours) 82.5 $\pm$6.96 81.8 $\pm$7.78 67.9 $\pm$5.43 72.25 $\pm$6.33
Increment in % 7.42 2.51 11.68 4.48
ScanObjectNN OBJ ONLY [SONN]
DCP-v1 [DCP] 51.7 $\pm$8.08 54.8 $\pm$8.82 32.15 $\pm$3.56 34.55 $\pm$3.17
MDCP-v1 (Ours) 54.7 $\pm$7.47 57.1 $\pm$6.51 39.0 $\pm$4.88 41.2 $\pm$5.72
Increment in % 5.80 4.20 21.31 19.25
DCP-v2 [DCP] 55.1 $\pm$9.57 56.2 $\pm$6.8 38.45 $\pm$5.08 42.05 $\pm$4.93
MDCP-v2 (Ours) 55.4 $\pm$12.11 62.4 $\pm$7.91 42.85 $\pm$3.37 46.15 $\pm$2.96
Increment in % 0.54 11.03 11.44 9.75
ScanObjectNN OBJ + BG [SONN]
DCP-v1 [DCP] 50.3 $\pm$5.25 48.2 $\pm$7.35 29.45 $\pm$5.34 32.15 $\pm$2.9
MDCP-v1 (Ours) 62.2 $\pm$6.73 60.2 $\pm$7.54 37.25 $\pm$3.26 41.45 $\pm$3.21
Increment in % 23.66 24.90 26.49 28.93
DCP-v2 [DCP] 52.3 $\pm$6.92 55 $\pm$8.16 33.8 $\pm$4.31 36.25 $\pm$4.57
MDCP-v2 (Ours) 58.8 $\pm$8.07 62.8 $\pm$6.22 41.9 $\pm$6.2 44.6 $\pm$4.1
Increment in % 12.43 14.18 23.96 23.03
ScanObjectNN PB75 [SONN]
DCP-v1 [DCP] 46.1 $\pm$8.31 45.2 $\pm$9.37 25.2 $\pm$3.78 28.45 $\pm$4.5
MDCP-v1 (Ours) 54.8 $\pm$7.35 51.5 $\pm$7.08 34.05 $\pm$4.2 37.65 $\pm$4.69
Increment in % 18.87 13.94 35.12 32.34
DCP-v2 [DCP] 49.6 $\pm$6.64 49.6 $\pm$8.7 27.85 $\pm$3.75 32.35 $\pm$4.03
MDCP-v2 (Ours) 54.7 $\pm$5.89 53.2 $\pm$9.46 34.45 $\pm$3.48 41.15 $\pm$4.87
Increment in % 10.28 7.26 23.70 27.20
Few-shot Classification. Table 2 evaluates the discriminative power of Mahalanobis distance. We compare features learned by DCP-v1/v2 [DCP] and MDCP-v1/v2 in few-shot classification tasks using the $k$-way, $m$-shot setting [CrossPoint]. Models are trained for registration on ModelNet40 [modelnet] and then tested on few-shot classification for ModelNet40 and three ScanObjectNN [SONN] splits (OBJ ONLY, OBJ+BG, PB75).
MDCP consistently outperforms DCP across all settings. Summarizing Table 2, MDCP-(v1,v2) achieves average accuracy improvement of $(22.98,6.52)$ on ModelNet40. For ScanObjectNN, improvements are: OBJ ONLY: $(12.68,8.19)$; OBJ+BG: $(25.99,18.40)$; PB75: $(25.06,17.11)$, [(a, b) denotes average accuracy increment (%) for MDCP-(v1,v2)].
These results suggest two key points: 1) v1 (without transformer) outperforms v2, indicating transformers may struggle with limited data; 2) Even on unseen ScanObjectNN data, MDCP shows consistent improvement across splits.
5.3 Endurance Test
Table 3: Quantitative comparison of MDCP-(v1,v2) robustness against original on ModelNet40 [modelnet] for low-density point cloud registration. We evaluate across various settings. MDCP consistently outperforms its original counterpart (bold denotes best). Lower is better for RMSE(t) (reported in $10^{-2}$ units). Note: -Geo refers to the vectorized Floyd-Warshall geodesic-graph version.
Target Low Density ModelNet40 [modelnet]
Unseen Data Unseen Category Noise
RMSE(R) RMSE(t) RMSE(R) RMSE(t) RMSE(R) RMSE(t)
1024 DCP-v1 [DCP] 42.0026 10.865 8.3181 1.2206 32.0181 0.9616
DCP-v1-Geo 38.1172 10.863 6.0088 1.2806 39.9996 0.9801
MDCP-v1 32.6880 10.977 5.8425 1.2096 24.9222 1.0062
DCP-v2 [DCP] 68.0209 3.6777 14.9977 2.7765 66.9599 3.0532
DCP-v2-Geo 44.6676 3.7786 14.7762 8.1678 48.2278 2.9987
MDCP-v2 34.9602 3.3871 14.3836 6.1023 25.8595 2.8308
512 DCP-v1 [DCP] 54.5579 1.6058 8.31497 1.6812 50.4602 1.4233
DCP-v1-Geo 48.1628 1.7898 8.1065 1.6711 49.7866 1.5661
MDCP-v1 42.3370 1.4918 8.4701 1.6028 38.0739 1.3168
DCP-v2 [DCP] 80.8099 8.9802 23.8370 3.7857 77.4393 8.7660
DCP-v2-Geo 77.8756 6.7752 21.8861 3.6544 48.2165 8.7765
MDCP-v2 47.3271 3.9172 28.7305 8.4551 44.3240 3.4637
Source Low Density ModelNet40 [modelnet]
1024 DCP-v1 [DCP] 43.2637 1.1146 8.304669 1.228 38.6104 1.0306
DCP-v1-Geo 36.6378 1.1123 8.0011 1.2761 39.1654 1.1756
MDCP-v1 31.3686 1.1035 8.3409 1.2247 26.9862 1.0158
DCP-v2 [DCP] 65.4777 3.1448 16.1993 1.9267 61.1659 2.6327
DCP-v2-Geo 57.1998 3.0098 14.7176 6.7861 49.1109 2.9987
MDCP-v2 38.6540 2.8759 7.4574 5.6750 27.1185 2.6114
512 DCP-v1 [DCP] 55.2426 1.5553 8.4311 1.4074 51.7135 1.3908
DCP-v1-Geo 38.1172 1.0863 6.0088 1.2806 39.9996 0.9801
MDCP-v1 42.4809 1.4700 7.8250 1.1089 39.6634 1.3272
DCP-v2 [DCP] 71.8377 7.2431 22.3028 4.8566 71.1976 6.9627
DCP-v2-Geo 44.6676 3.7698 26.8861 4.9908 65.9981 5.0098
MDCP-v2 50.1560 3.6758 29.2851 4.4566 45.7104 3.3395
Efficiency towards Point-density. Table 3 evaluates MDCP’s robustness against varying point densities, simulating real-world sensor data [survey]. One point cloud (source or target) is downsampled by half (e.g., 2048 → 1024). Evaluation metrics use the same number of points for both clouds (e.g., source with 1024 points and target with 2048 points are used to compute transformation, then applied to original points for error calculation).
Figure 3 compares DCP-v1 [DCP] and MDCP-v1. When densities are similar (man’s face), both perform well. However, MDCP-v1 demonstrates superior robustness under varying densities (chair). This highlights its effectiveness in real-world applications where source or target point clouds may have different densities due to factors like sensor capabilities or data acquisition methods (e.g., SFM-LiDAR fusion).
Note: Source point cloud is a transformed version of $g(Y)$, and the goal is to estimate $g$ for perfect alignment with target $X$.
Table 4: Quantitative comparison of proposed MDeepUME’s robustness on ModelNet40 [modelnet], FAUST [FAUST], and Stanford 3D datasets. We evaluate across various noise, benchmarked by authors in DeepUME [DeepUME]. MDeepUME consistently outperforms DeepUME (bold denotes best results). Lower is better for all metrics: RMSE(t), chamfer distance (CD) [ChamferDistance], and Hausdorff distance (HD) [HausdorffDistance] (reported in $10^{-2}$ units).
Noise-types DeepUME [DeepUME] MDeepUME
CD HD RMSE(R) RMSE(t) CD HD RMSE(R) RMSE(t)
MN 40 Bernoulli 1.113 8.6300 46.39461 1.5381 1.1120 8.6200 45.7815 1.5398
Gaussian 0.1913 1.2445 2.526296 0.0625 0.1900 1.2400 1.9946 0.0642
Sampling 0.6317 5.6678 31.39413 0.9197 0.6290 5.6620 30.7709 0.9236
Z-Intersection 1.2237 11.3676 89.48223 0.9003 1.2230 11.3670 87.4486 0.911
FAUST Bernoulli 0.2608 2.7165 11.30384 2.2294 0.2728 2.7818 10.3421 2.1964
Gaussian 0.1243 0.9342 1.819818 0.1085 0.1230 0.9300 1.7029 0.2692
Sampling 0.1026 0.9342 5.6842 1.172 0.0941 0.9320 3.7985 0.9461
Z-Intersection 0.2064 2.3368 12.3721 2.2183 0.2000 2.3350 11.9663 2.1383
S-3D Bernoulli 0.2345 10.4470 5.951346 1.2268 0.2470 10.1853 6.4044 1.2266
Gaussian 0.1068 0.9477 0.392542 0.0591 0.1040 0.9410 0.3106 0.0604
Sampling 0.0908 8.5766 6.709815 0.7145 0.0900 8.4267 6.1997 0.7143
Z-Intersection 0.1771 11.7627 5.66907 1.0417 0.1700 11.5800 4.4850 1.0428
Efficiency towards Various Noise Types. Table 4 evaluates the robustness of MDeepUME against various noise types from DeepUME [DeepUME] (Bernoulli, Gaussian, Sampling, Zero-Intersection). We benchmark on ModelNet40 [modelnet] (training and evaluation), FAUST [FAUST], and Stanford3D (testing only). Training settings follow the original paper [DeepUME] ( [^3] https://github.com/langnatalie/DeepUME).
MDeepUME outperforms DeepUME across datasets and noise types in terms of Chamfer distance (CD) [ChamferDistance], Hausdorff distance (HD) [HausdorffDistance], and root mean squared error in rotation (RMSE(R)). However, limitations are observed in translation accuracy (RMSE(t)) due to DeepUME’s inherent projection into SO(3)-invariant space, where Mahalanobis distance is less effective.
6 Conclusion
Our paper introduces Mahalanobis k-NN, a statistical framework for improving point cloud registration in learning-based methods. It addresses challenges like feature matching due to variations in point cloud densities. Mahalanobis k-NN leverages local neighborhood distribution for accurate feature extraction, outperforming methods like Flyod-Warshall. It integrates seamlessly into local graph-based point cloud analysis methods like DCP and DeepUME, achieving state-of-the-art performance on benchmark datasets. We also demonstrate that registered point cloud features possess discriminative capabilities, leading to significant improvements in few-shot classification tasks. Mahalanobis k-NN offers exceptional surface awareness and outlier robustness, benefiting tasks like segmentation, decomposition, and normal estimation. Additionally, it serves as a surficial geometry-based error metric, enhancing tasks like generation and upsampling. Its versatility makes it a crucial component in advancing point cloud processing, warranting further benchmarking to highlight its significance.
[…]

### v2

@@
 Abstract
-In this paper, we discuss Mahalanobis k-NN: a statistical lens designed to address the challenges of feature matching in learning-based point cloud registration when confronted with an arbitrary density of point clouds, either in the source or target point cloud.
+In this paper, we discuss Mahalanobis k-NN: A Statistical Lens designed to address the challenges of feature matching in learning-based point cloud registration when confronted with an arbitrary density of point clouds.
 We tackle this by adopting Mahalanobis k-NN’s inherent property to capture the distribution of the local neighborhood and surficial geometry.
@@
 Moreover, we establish for the first time that the features acquired through point cloud registration inherently can possess discriminative capabilities.
-This is evident by a substantial improvement of about 20% in the average accuracy observed in the point cloud few-shot classification task benchmarked on ModelNet40 and ScanObjectNN.
-The code is publicly available at https://github.com/TejasAnvekar/Mahalanobis-k-NN
-Keywords:
-Point Cloud Registration Metric Distance 3D Vision.
-Figure 1: The visual supremacy of our proposed methodologies, MDCP-v1 and MDeepUME, becomes apparent in a point cloud registration task involving a target point cloud with only half the points compared to the source point cloud.
-In the visualization, highlighted boxes in red illustrate the limited performance of DCP-v1 [DCP] and DeepUME [DeepUME], while the highlighted boxes in green demonstrate the resilience of the proposed Mahalanobis versions of DCP-v1 [DCP] and DeepUME [DeepUME].
+This is evident by a substantial improvement of about 20% in the average accuracy observed in the point cloud few-shot classification task, benchmarked on ModelNet40 and ScanObjectNN.
+Figure 1: The visual supremacy of our proposed methodologies, MDCP-v1 and MDeepUME, becomes apparent in a point cloud registration, involving a target point cloud with only half the points compared to the source point cloud.
+In the visualization, regions highlighted in red illustrate the limited performance of DCP-v1[42] and DeepUME[27], while the regions highlighted in green demonstrate the resilience of the proposed Mahalanobis versions of DCP-v1[42] and DeepUME[27].
 1 Introduction
-With the rapid progress of high-precision sensors like LiDAR [lidar] and Kinect [kinect], point clouds have become the prevalent data format for 3D representation.
+With the rapid progress of high-precision sensors like LiDAR [13] and Kinect [52], point clouds are prevalent data format for 3D representation.
 Due to sensor limitations in capturing only partial views, registration algorithms are crucial for amalgamating scans into comprehensive 3D scenes.
 Point cloud registration involves determining transformation matrices between point cloud pairs, facilitating the fusion of partial scans to produce coherent 3D representations.
-Registration is a fundamental task spanning various computational fields such as medical imaging [medicalImaging], robotics [robotics], autonomous driving [autoDriving], and computational chemistry [ComChe].
-Estimating transformation matrices serves applications like 3D reconstruction [reconstruction] and precise 3D localization [Localization].
-It underpins the development of intricate 3D maps for autonomous driving [autoDriving], environment reconstruction in robotics [robotics], and improved safety in mining.
-Additionally, registration’s ability to facilitate high-precision localization proves invaluable for entities like driverless cars [autoDriving], ensuring precise positioning and interaction within the 3D environment.
-The registration of point clouds has garnered significant attention through closed-form [ICP, goICP, FGR, aligningPointCloudView, ume] and learning-based [pointnetLK, DCP, DeepGMR, DeepUME, rgm] methodologies.
-DCP [DCP] tackles the challenge of feature matching (correspondences) by initially estimating local features via weight-shared DGCNN [dgcnn] (Dynamic Graph Convolutional Neural Network).
+Registration is a fundamental task spanning various computational fields such as medical imaging [9], robotics [32], autonomous driving [12], and computational chemistry [25].
+Estimating transformation matrices serves applications like 3D reconstruction [21] and precise 3D localization [22].
+It underpins the development of intricate 3D maps for autonomous driving [12], environment reconstruction in robotics [32], and improved safety in mining.
+Additionally, registration’s ability to facilitate high-precision localization proves invaluable for entities like driverless cars [12], ensuring precise positioning and interaction within the 3D environment.
+The registration of point clouds has garnered significant attention through closed-form [4, 47, 53, 37, 15] and learning-based [2, 42, 50, 27, 18] methodologies.
+DCP [42] tackles the challenge of feature matching (correspondences) by initially estimating local features via weight-shared DGCNN [44] (Dynamic Graph Convolutional Neural Network).
 Subsequent steps involve an attention-based module and differentiable SVD modules for point-to-point registration.
-Conversely, DeepUME [DeepUME] adopts a distinct approach by projecting points into a space invariant under $\mathcal{SO}(3)$ transformations.
+Conversely, DeepUME [27] adopts a distinct approach by projecting points into a space invariant under $\mathcal{SO}(3)$ transformations.
 This space is leveraged to compute invariant per-point features that address real-world data sampling issues.
-Employing two weight-shared transformers and DGCNN [dgcnn], correspondences are established and fed into UME [ume] for rigid transformation parameter estimation.
-A limitation of both DCP [DCP] and DeepUME [DeepUME] is susceptibility to failure in scenarios where either the source or target point cloud exhibits arbitrary (lower) density than the others—an occurrence commonly encountered in 3D scan registration from diverse sensors.
-This limitation is attributed to the methods’ reliance on estimating per-point correspondences using local features extracted from edge-conv [dgcnn], which operates on graphs constructed using k-nearest neighbors (k-NN).
-We introduce Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration to mitigate this issue.
-The efficacy of Mahalanobis distance as an evaluation metric has already been demonstrated in reference-based point cloud quality assessment [MD1].
+Employing two weight-shared transformers and DGCNN [44], correspondences are established and fed into UME [15] for rigid transformation parameter estimation.
+A limitation of both DCP [42] and DeepUME [27] is susceptibility to failure in scenarios where either the source or target point cloud exhibits arbitrary (lower) density than the others—an occurrence commonly encountered in 3D scan registration from diverse sensors.
+This limitation is attributed to the methods’ reliance on estimating per-point correspondences using local features extracted from edge-conv [44], which operates on graphs constructed using k-nearest neighbors (k-NN).
+We introduce Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration to mitigate aforementioned issue.
+The efficacy of Mahalanobis distance as an evaluation metric has already been demonstrated in reference-based point cloud quality assessment [24].
 Our proposed Mahalanobis k-NN can be used as a plugin for any point-cloud registration method.
-In this paper, we propose mahalnobis versions of DCP-(v1,v2) [DCP] and DeepUME [DeepUME].
+In this paper, we propose mahalnobis versions of DCP-(v1,v2) [42] and DeepUME [27].
 Both methods operate on diverse point-cloud representation spaces.
@@
 Additionally, we demonstrate the discriminative capabilities of the proposed Mahalanobis lens through point cloud few-shot classification tasks.
-In this context, models pre-trained for registration tasks leverage DGCNN [dgcnn] features for few-shot evaluation on both ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets.
+In this context, models pre-trained for registration tasks leverage DGCNN [44] features for few-shot evaluation on both ModelNet40 [45] and ScanObjectNN [41] datasets.
 Finally, we summarize our contributions as follows:
   • We propose Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration by incorporating Mahalanobis distance in:
-    • Deep Closest Point (DCP) [DCP] that operates on Euclidean coordinate space.
-    • Deep Universal Manifold Embedding (DeepUME) [DeepUME] that operates on $\mathcal{SO}(3)$ invarient space.
+    • Deep Closest Point (DCP) [42] that operates on Euclidean coordinate space.
+    • Deep Universal Manifold Embedding (DeepUME) [27] that operates on $\mathcal{SO}(3)$ invariant space.
   • We demonstrate the efficacy of the proposed method on various point cloud registration tasks on a variety of publicly available data sets and compare the results with state-of-the-art techniques.
@@
 We achieve state-of-the-art results compared to other point cloud registration methods on all benchmarking strategies and on almost all evaluation metrics.
-  • To the best of our knowledge, we are the first to demonstrate discriminative prowess inherited in point cloud registration task by benchmarking point cloud few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets through features extracted by DCP [DCP], we report up to an average of 20% increment when Mahalanobis is incorporated.
-  • We will release our code to facilitate reproducibility and future research upon acceptance.
+  • To the best of our knowledge, we are the first to demonstrate discriminative prowess inherited in point cloud registration task by benchmarking point cloud few-shot classification on ModelNet40 [45] and ScanObjectNN [41] datasets through features extracted by DCP [42], we report up to an average of 20% increment when Mahalanobis is incorporated.
 2 Related Works
 Optimization Based.
-Preceding the advent of the deep learning era in 3D point cloud registration, one common strategy involved extracting and matching spatially local features, as demonstrated in studies like [comprehensiveEvalFeatureDescriptor, spinImages, aligningPointCloudView, RepeatabilityRobust3Dmatching, fastRobustDescriptor, rotationalContour].
-Many existing methods in this category are adaptations of 2D image processing solutions, such as variants of 3D-SIFT [3Dsift] and the 3D Harris key-point detector [harris3D].
+Preceding the advent of the deep learning era in 3D point cloud registration, one common strategy involved extracting and matching spatially local features, as demonstrated in studies like [19, 26, 37, 48, 46, 49].
+Many existing methods in this category are adaptations of 2D image processing solutions, such as variants of 3D-SIFT [29] and the 3D Harris key-point detector [40].
 However, key-point matching in 3D presents challenges due to the absence of a regular sampling grid, artifacts, and sampling noise, leading to high outlier rates and localization errors.
-To achieve global alignment, key-point matching typically utilizes outlier rejection methods like RANSAC [ransac] and is refined using local optimization algorithms [ICP, scanRegistration3Dndt, registrationWithoutICP, iterativePointMatching].
-Notably, DGR [DGR] follows a similar paradigm, but it incorporates learnable inlier detection.
-Researchers have proposed numerous works for handling outliers and noise [trimmedICP], formulating robust minimizers [robustRegistration2D3D], and devising more suitable distance metrics.
-The widely used Iterative Closest Point algorithm (ICP) [ICP, iterativePointMatching], a popular refinement algorithm, constructs point correspondences based on spatial proximity and employs a transformation estimation step.
-Over time, various ICP variants [efficientICP, goICP, sparseICP] have been proposed to enhance convergence rate, robustness, and accuracy in the 3D point cloud registration.
+To achieve global alignment, key-point matching typically utilizes outlier rejection methods like RANSAC [16] and is refined using local optimization algorithms [4, 30, 33, 51].
+Notably, DGR [11] follows a similar paradigm, but it incorporates learnable inlier detection.
+Researchers have proposed numerous works for handling outliers and noise [10], formulating robust minimizers [17], and devising more suitable distance metrics.
+The widely used Iterative Closest Point algorithm (ICP) [4, 51], a popular refinement algorithm, constructs point correspondences based on spatial proximity and employs a transformation estimation step.
+Over time, various ICP variants [36, 47, 8] have been proposed to enhance convergence rate, robustness, and accuracy in the 3D point cloud registration.
 Learning Based.
-Registration approaches leverage learning-based techniques alongside closed-form methods.
-Pioneered by PointNet [pointnet] and advanced by DGCNN [dgcnn], learning from data for point cloud registration has become powerful (e.g., PointNet++, DCP [DCP], PRNet [prnet]).
-PointNetLK [pointnetLK] minimizes learned feature distance using a differentiable Lucas-Kanade algorithm, while DCP uses attention and differentiable SVD for point-to-point registration.
-Recent work [rgm] introduces graph transformation and deep graph matching.
-Moreover, DeepUME [DeepUME] employ $\mathcal{SO}(3)$ invariant coordinate systems for global registration.
-DeepGMR [DeepGMR] tackles pose-invariant correspondences between raw point clouds and Gaussian-mixture models but suffers from noise sensitivity.
-DeepUME [DeepUME] fuses closed-form Universal Manifold Embedding (UME) and deep learning to handle sparse and unevenly sampled point clouds with large transformations.
-This framework is trained end-to-end and learns joint-resampling strategies and invariant features.
-DCP [DCP] and DeepUME [DeepUME] achieve impressive performance but struggle when the source or target point cloud densities are unequal.
-Both rely on DGCNN [dgcnn] for local feature-based correspondence estimation.
-However, DGCNN’s Euclidean metric for k-nearest neighbors lacks surface awareness.
-We propose Mahalanobis versions of DCP and DeepUME, introducing the Mahalanobis distance for surface-aware features and enhanced registration in real-world scenarios.
+Apart from optimization-based methods, registration approaches have also leveraged learning-based techniques.
+Pioneered by PointNet  [34] and further developed by DGCNN [44], learning from data in a task-specific manner has become a powerful tool in point cloud registration.
+These learned point cloud representations can be exploited for robust registration, as evidenced in studies like [35, 42, 43].
+For instance, PointNetLK  [2] minimizes learned feature distance using a differentiable Lucas-Kanade algorithm  [28], while DCP [42] addresses feature matching through attention-based modules and differentiable SVD modules for point-to-point registration.
+Recently, [18] introduces a transformation of point clouds into graphs and utilizes deep graph matching to extract deep feature soft correspondence matrices.
+To overcome this limitation, researchers have proposed approaches [27] employing $\mathcal{SO}(3)$ invariant coordinate systems to provide a global registration solution.
+Another method, DeepGMR  [50], deals with pose-invariant correspondences between raw point clouds and Gaussian mixture model (GMM) parameters, recovering transformations from the matched mixtures, although it exhibits performance degradation in the presence of sampling noise.
+Moreover, DeepUME [27] lies in its innovative fusion of the closed-form Universal Manifold Embedding (UME) method and a deep neural network to address the challenge of registering sparsely and unevenly sampled point clouds subjected to large transformations.
+This unified framework, trained end-to-end and without supervision, employs an $\mathcal{SO}(3)$-invariant coordinate system to learn joint-resampling strategies and invariant features.
+While DCP [42] and DeepUME [27] exhibit impressive performance across various real-world scenarios in point cloud registration, such as robustness to diverse noise types, they face limitations when dealing with imbalanced point cloud densities.
+Both methods rely on correspondence estimation based on local features using DGCNN [44].
+However, DGCNN’s intrinsic lack of surface awareness due to its utilization of the Euclidean metric for k-nearest neighbors introduces challenges [38, 39].
+To overcome this, authors in [38, 39] introduced using ellipsoid query but falls short as it requires extraneous querying.
+Counter to this we propose Mahalanobis versions of both DCP and DeepUME, introducing the Mahalanobis distance into k-NN to derive surface-aware features that enhance point cloud registration.
 Furthermore, our proposed approach demonstrates remarkable discriminative capabilities, as evident from its performance in point cloud few-shot classification tasks.
@@
 The term $d\big{(}X,g(Y)\big{)}$ is a measure of the projection error between $X$ and the transformed $Y$, expressed as $d\big{(}X,RY+t\big{)}=\sum_{s=1}^{S}\left\|x_{s}-\big{(}Ry_{s}+t\big{)}\right\|_{2}$, where $s$ iterates over the $S$ pairs of corresponding points.
-The optimization problem in Equation 1 embodies a well-known “chicken-and-egg” scenario: determining the optimal transformation matrix requires knowledge of true correspondences  [correspondance1, correspondance2]; conversely, accurate correspondences can be established given the optimal transformation matrix.
+The optimization problem in Equation 1 embodies a well-known “chicken-and-egg” scenario: determining the optimal transformation matrix requires knowledge of true correspondences  [5, 6]; conversely, accurate correspondences can be established given the optimal transformation matrix.
 However, the simultaneous resolution of both aspects poses a non-trivial challenge.
 4 Method
-In this paper, we focus on point cloud registration using DCP [DCP] and DeepUME [DeepUME] for global one-shot rigid transformations (contrasting iterative methods like ICP [ICP]).
-Both rely on DGCNN [dgcnn] for local features but use static k-nearest neighbor (k-NN) graphs, limiting semantic space creation.
-Unlike, DGCNN’s dynamic k-NN graphs, these methods restrict feature projection depth.
-We propose using the Mahalanobis distance [MD] to enhance k-NN, enabling surface-aware neighbor selection based on principal components.
-This extracts surface and corner features crucial for accurate matching and robust transformation estimation.
+In this paper, we focus on two deep learning-based approaches, specifically DCP  [42] and DeepUME  [27], which address the aforementioned challenge as a global one-shot rigid transformation.
+This contrasts with the iterative nature of methods like ICP  [4].
+Considering these methods, DGCNN  [44] is pivotal for extracting local features crucial for high-quality point matching.
+Both approaches modify the DGCNN framework to construct a static graph using k-nearest Neighbors (k-NN) for the point cloud.
+However, this limits the potential for complex semantic space creation.
+The conventional DGCNN recalculates the k-NN graph after each convolution, projecting features into higher-dimensional space, while the simplified k-NN graph imposition curtails semantic depth.
+To overcome this limitation, we propose harnessing the Mahalanobis distance [31] as a statistical lens for enhancing k-NN.
+This empowers querying based on principal components, emphasizing surface-level k-NN grouping.
+The Mahalanobis-based k-NN lens extracts features sensitive to surface and corner regions, vital for accurate feature matching and robust transformation estimation.
 4.1 Mahalanobis k-NN
-Figure 2: The illustration contrasts Euclidean distance fields (red) capturing spatial neighbors with Mahalanobis distance fields (blue) considering the underlying data distribution.
-A chair point cloud exemplifies this.
-The black query point is surrounded by Euclidean neighbors (red points) and Mahalanobis neighbors (blue points).
-The depiction clearly illustrates the impact of Mahalanobis distance, effectively capturing surface points per the data distribution vital for precise feature matching.
-In this section, we introduce Mahalanobis distance as a statistical lens for robust feature matching in DCP [DCP] and DeepUME [DeepUME].
+Figure 2: Illustration highlights the distinction between the Red gradients representing Euclidean distance fields that capture spatial neighbors and the Blue gradients depicting Mahalanobis distance fields, which consider neighbors concerning the underlying data distribution.
+To further emphasize the point, we present the influence of both Euclidean and Mahalanobis k-NN on a chair point cloud.
+The black query point is surrounded by Euclidean neighbors shown in red points and Mahalanobis neighbors in blue points.
+The depiction clearly illustrates the impact of Mahalanobis distance, effectively capturing surficial points per the data distribution—vital for precise feature matching.
+In this section, we introduce Mahalanobis distance as a statistical lens for robust feature matching in DCP [42] and DeepUME [27].
 We modify DGCNN to construct a graph $\mathcal{G}^{\dagger}$ using Mahalanobis k-nearest neighbors (k-NN) for enhanced feature selection.
@@
 where $\mathcal{N}_{i}$ denotes neighbors of vertex $i$ in $\mathcal{G}$.
-We propose a novel augmentation by integrating Mahalanobis k-nearest neighbors (k-NN) in Equation.2 to construct graph $\mathcal{G}^{\dagger}$.
+We propose a novel augmentation by integrating Mahalanobis k-nearest neighbors (k-NN) in Equation 2 to construct graph $\mathcal{G}^{\dagger}$.
 This graph prioritizes surface awareness through selection of surficial neighbors $\mathcal{N}_{i}^{\dagger}$ based on Mahalanobis distance as shown in Figure 2.
@@
 To counter this, we implement vectorized Floyd-Warshall (refer supplementary for implementation detail on vectorized Floyd-Warshall) on the Euclidean-knn graph and observe that the performance is slower and also not as effective as the proposed Mahalanobis variants of DCP and DeepUME.
+Advantages of Mahalanobis Distance.
+Unlike Euclidean distance which treats dimensions independently, Mahalanobis distance accounts for covariance, making it suitable for correlated feature datasets.
+It scales features invariantly, accommodating feature scale variations that affect Euclidean distances.
+This metric excels for datasets with non-spherical or elliptical clusters, considering data distribution shape.
+As data dimensionality increases, Mahalanobis distance is more resilient against the “curse of dimensionality,” particularly impactful on Euclidean distances.
+It demonstrates robustness to outliers, focusing on central data distribution.
+Moreover, Mahalanobis distance enables customized metric learning, adapting to dataset peculiarities for improved performance.
+In machine learning, it serves as a discriminative metric, enhancing class separation in classification.
+Its statistical interpretation enhances usability in data analysis.
+Moreover, it’s faster and more robust compared to geodesic-graph-based approaches.
 The impact of Mahalanobis distance is visually depicted in Figure 2.
 This surficial awareness proves crucial in feature matching.
-Incorporating Mahalanobis distance into DCP  [DCP] and DeepUME  [DeepUME] yields superior performance in point cloud registration.
+Incorporating Mahalanobis distance into DCP  [42] and DeepUME  [27] yields superior performance in point cloud registration.
 Discriminative power is affirmed in point-cloud few-shot classification, with pre-trained DGCNN models from vanilla DCP and Mahalanobis DCP showcasing remarkable classification capabilities.
 5 Experiments
-In this section, we showcase the superior performance of our proposed method in the domain of point cloud registration when compared against existing state-of-the-art techniques such as ICP [ICP], GO-ICP [goICP], FGR [FGR], PointNetLK [pointnetLK], DCP [DCP], and DeepUME [DeepUME] on publicly available datasets.
+In this section, we showcase the superior performance of our proposed method in the domain of point cloud registration when compared against existing state-of-the-art techniques such as ICP [4], GO-ICP [47], FGR [53], PointNetLK [2], DCP [42], and DeepUME [27] on publicly available datasets.
 Our evaluation encompasses a spectrum of benchmarking strategies, encompassing scenarios where one point cloud is notably sparser than the other.
@@
 5.1 Point Cloud Registration
-We build upon the model architecture, training strategies, evaluation metrics, and hyperparameters established in DCP [DCP]  [^1] https://github.com/WangYueFt/dcp.
+We build upon the model architecture, training strategies, evaluation metrics, and hyperparameters established in DCP [42]  [^1] https://github.com/WangYueFt/dcp.
 We introduce the Mahalanobis version of DCP (MDCP), detailed in Section 4.1.
 MDCP comes in two variants: v1 without a transformer and v2 with a transformer.
-We further extend our approach to DeepUME [DeepUME] by proposing the Mahalanobis variant, MDeepUME.
+We further extend our approach to DeepUME [27] by proposing the Mahalanobis variant, MDeepUME.
 For a fair comparison, we benchmark geodesic versions of these models using vectorized Floyd-Warshall on Euclidean k-NN graphs.
 Dataset.
-To facilitate benchmarking, we employ publicly available datasets: ModelNet40 [modelnet], FAUST [FAUST], and Stanford3D (S-3D) [^2] http:// graphics.stanford.edu/data/3Dscanrep/ .
+To facilitate benchmarking, we employ publicly available datasets: ModelNet40 [45], FAUST [7], and Stanford3D (S-3D) [^2] http:// graphics.stanford.edu/data/3Dscanrep/ .
 Notably, the latter two datasets are exclusively employed for testing purposes.
-To ensure methodological consistency, we adhere to the official training/testing splits and settings as stipulated in the original works of DCP [DCP] and DeepUME [DeepUME].
+To ensure methodological consistency, we adhere to the official training/testing splits and settings as stipulated in the original works of DCP [42] and DeepUME [27].
 Among these datasets, our approach is rigorously evaluated.
@@
 5.2 Comparison with State-of-the-art Methods
-Table 1: Comparison on ModelNet40 [modelnet] for unseen data, unseen category, and Gaussian noise.
-MDCP-(v1,v2) and MDeepUME outperform baselines (bold best, underline second best).
-Lower is better.
-$\dagger$ denotes reproduced results, -Geo refers to Floyd-Warshall geodesic-graph version.
+Table 1: Quantitative comparison of state-of-the-art methods benchmarked on ModelNet40 [45] Dataset for three tasks 1) Unseen data, 2) Unseen category, 3) Robustness towards Gaussian Noise; Our proposed MDCP-(v1,v2) and MDeepUME; outperforms its original variants in almost all cases.
+All evaluation metrics are lower the better.
+We highlight bold as best and underline as second best.
+Note: $\dagger$ depicts results reproduced on our systems and -Geo refers to Flyod-Warshall geodesic-graph version.
 Unseen data Unseen Category Unseen data + Noise
 RMSE(R) RMSE(t) RMSE(R) RMSE(t) RMSE(R) RMSE(t)
-ICP [ICP] 29.9148 0.2909 29.8764 0.2933 29.7080 0.2906
-GO-ICP [goICP] 11.8523 0.0257 13.8657 0.0226 11.4535 0.0231
-FGR [FGR] 9.3628 0.0139 9.8490 0.0135 24.6515 0.1090
-PointNetLK [pointnetLK] 15.0954 0.0221 17.5021 0.0280 16.0049 0.0216
-[…]
-[…]
-[…]
+ICP [4] 29.9148 0.2909 29.8764 0.2933 29.7080 0.2906
+GO-ICP [47] 11.8523 0.0257 13.8657 0.0226 11.4535 0.0231
+FGR [53] 9.3628 0.0139 9.8490 0.0135 24.6515 0.1090
+PointNetLK [2] 15.0954 0.0221 17.5021 0.0280 16.0049 0.0216
+DCP-v1 [42] 2.5457 0.0018 4.3819 0.0050 2.6318 0.0018
+DCP-v1 [42] 2.6581 0.0009 6.8386 0.0035 2.6048 0.0009
+DCP-v1-Geo 2.6001 0.0007 3.3326 0.0017 2.1778 0.0019
 MDCP-v1 (Ours) 1.6614 0.0004 1.9679 0.0004 1.5889 0.0004
-[…]
-[…]
-[…]
+DCP-v2 [42] 1.1434 0.0018 3.1502 0.0050 1.0814 0.0015
+DCP-v2 [42] 1.4125 0.0018 4.2246 0.0063 1.1479 0.0016
+DCP-v2-Geo 1.4478 0.0017 5.5233 0.0051 2.1668 0.0019
 MDCP-v2 (Ours) 0.9468 0.0004 1.3849 0.0049 0.8481 0.0043
-DeepUME [DeepUME] 0.0062 0 0.0183 0.00019 2.5263 0.0006
-[…]
+DeepUME [27] 0.0062 0 0.0183 0.00019 2.5263 0.0006
+DeepUME-Geo 0.0051 0 0.0106 0.00014 2.0086 0.0008
 MDeepUME 0.0023 0 0.0098 0.00009 1.9946 0.0006
-ModelNet40 Results.
-Table 1 compare our methods (MDCP, MDeepUME) with baselines and their geodesic versions.
-We outperform all methods across various settings: Unseen data: Train-test split from ModelNet40 [modelnet]; Unseen category: First 20 classes for training, remaining for testing; Gaussian Noise: $\mathcal{N}(0,0.01)$ clipped to [-0.05, 0.05].
-The point density is 1024 in all settings.
-Baseline evaluations are from [DCP] (highly comparable results achieved with minimal standard deviation during our reproduction attempts).
-For each case, a rigid transformation is applied along each axis (rotation: $[0^{\circ},45^{\circ}]$, translation: [-0.5, 0.5]).
-The network receives both the original and transformed point cloud as input, evaluated against ground truth using MDCP and MDeepUME.
-Figure 3: Robustness evaluation on unseen data (FAUST [FAUST] hand), low-density source (airplane), and low-density target (chair) point clouds (trained on ModelNet40 [modelnet]).
-Red boxes highlights vulnerabilities in DCP-v1 [DCP], while green boxes highlights MDCP-v1’s efficacy.
-Table 2: Evaluation of few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] (OBJ ONLY, OBJ+BG, PB75 splits).
-Our DCGNN in MDCP-(v1,v2) outperforms baselines across all settings (bold denotes best accuracy).
-We report mean accuracy (%) and standard deviation from 50 runs.
-ModelNet40 [modelnet]
+Results on ModelNet40 Dataset [45].
+Comprehensive results are presented in Table 1, offering a thorough comparison between our method, state-of-the-art alternatives and their geodesic versions.
+The presented outcomes highlight the superior performance of our method across all evaluation metrics and in diverse settings.
+Our benchmarking encompasses scenarios including Unseen-data evaluation, where the train-test split outlined in ModelNet40 [45] is utilized; Unseen category evaluation, where the first 20 classes of ModelNet40 are employed for training and the remaining for testing; and a robustness assessment against Gaussian Noise $\mathcal{N}(0,0.01)$, which is clipped to the interval [-$0.05,0.05$] before being added to the point cloud.
+Notably, a point density of 1024 is consistently sampled in all aforementioned settings.
+Note: Our assessment of state-of-the-art methods is derived from [42], as our efforts to reproduce the open-source code yielded highly comparable results with minimal standard deviation.
+In each case, a rigid transformation is introduced along each axis, with rotation uniformly sampled in the range [$0^{\circ},45^{\circ}$] and translation within [-$0.5,0.5$].
+The input to the network consists of both the original point cloud and the transformed point cloud through the rigid motion.
+This input is subsequently evaluated against the known ground truth in both our proposed MDCP and MDeepUME methods.
+Figure 3: Evaluation of Robustness for varied test cases.
+Both DCP-v1 [42] and proposed MDCP-v1 trained on ModelNet40 [45] dataset.
+Our evaluation encompasses three distinct scenarios: 1) Evaluation on unseen data, specifically a man holding palms together from the FAUST [7] dataset; 2) Assessment of low-density source point clouds, exemplified by an airplane; and 3) Examination of low-density target point clouds, illustrated by a chair.
+In our visual analysis, regions highlighted in red highlights the vulnerabilities observed in vanilla DCP-v1 [42], while regions in green accentuate the pronounced efficacy demonstrated by the proposed MDCP-v1 over the conventional DCP-v1.
+Figure 4: We present a compelling demonstrations showcasing the surface awareness within the proposed MDCP-v1, in contrast to the Euclidean approach in DCP-v1[42].
+Each row within the illustration corresponds to K values ranging from 2 to 5, effectively exemplifying the pronounced surficial awareness embodied by MDCP-v1.
+This surficial awareness is pivotal for robust point cloud registration tasks.
+Furthermore, to substantiate the assertion that Mahalanobis distance operates as a statistical lens for enhanced feature matching, we undertake KMeans [20] clustering within the initial embedding space of DGCNN.
+This embedding space, a 64-dimensional domain following the first edge-convolution [44], is employed both in DCP-v1 and MDCP-v1.
+Our findings are presented visually through Figure 4.
+The depicted results unequivocally showcase the preservation of surface awareness in the case of the Mahalanobis variant, even in scenarios involving overclustering (e.g., K=5).
+This outcome stands in contrast to its Euclidean counterpart.
+Particularly is the consistent maintenance of surface awareness throughout various clustering cases, particularly evident in the context of the Airplane dataset.
+This affirmation highlights that the Mahalanobis version of DCP learns to discern and emphasize surface-aware regions that are crucial in achieving precise feature matching.
+Table 2: We present the outcomes of few-shot classification experiments conducted on the ModelNet40 [45] dataset, as well as on three distinct splits (OBJ ONLY, OBJ+BG, and PB75) of the ScanObjectNN [41] dataset.
+Our proposed DCGNN employed in MDCP-(v1,v2) consistently surpasses its original variants by a substantial margin across all few-shot settings and datasets.
+The reported quantitative results in terms of accuracies (%) are presented below, with values in bold indicating the best performance.
+Additionally, we provide the standard deviation values derived from 50 runs for a comprehensive understanding of the results.
+ModelNet40 [45]
 5 way 10 way
 Method 10 shot 20 shot 10 shot 20 shot
-DCP-v1 [DCP] 67.6 $\pm$6.91 69 $\pm$7.87 50.5 $\pm$7.74 57.1 $\pm$5.01
+DCP-v1 [42] 67.6 $\pm$6.91 69 $\pm$7.87 50.5 $\pm$7.74 57.1 $\pm$5.01
 MDCP-v1 (Ours) 75.2 $\pm$7.3 80 $\pm$7.74 71.0 $\pm$4.16 70.9 $\pm$5.72
 Increment in % 11.24 15.94 40.59 24.17
-DCP-v2 [DCP] 76.8 $\pm$4.01 79.8 $\pm$6.17 60.8 $\pm$4.93 69.15 $\pm$5.7
+DCP-v2 [42] 76.8 $\pm$4.01 79.8 $\pm$6.17 60.8 $\pm$4.93 69.15 $\pm$5.7
 MDCP-v2 (Ours) 82.5 $\pm$6.96 81.8 $\pm$7.78 67.9 $\pm$5.43 72.25 $\pm$6.33
 Increment in % 7.42 2.51 11.68 4.48
-ScanObjectNN OBJ ONLY [SONN]
-DCP-v1 [DCP] 51.7 $\pm$8.08 54.8 $\pm$8.82 32.15 $\pm$3.56 34.55 $\pm$3.17
+ScanObjectNN OBJ ONLY [41]
+DCP-v1 [42] 51.7 $\pm$8.08 54.8 $\pm$8.82 32.15 $\pm$3.56 34.55 $\pm$3.17
 MDCP-v1 (Ours) 54.7 $\pm$7.47 57.1 $\pm$6.51 39.0 $\pm$4.88 41.2 $\pm$5.72
 Increment in % 5.80 4.20 21.31 19.25
-DCP-v2 [DCP] 55.1 $\pm$9.57 56.2 $\pm$6.8 38.45 $\pm$5.08 42.05 $\pm$4.93
+DCP-v2 [42] 55.1 $\pm$9.57 56.2 $\pm$6.8 38.45 $\pm$5.08 42.05 $\pm$4.93
 MDCP-v2 (Ours) 55.4 $\pm$12.11 62.4 $\pm$7.91 42.85 $\pm$3.37 46.15 $\pm$2.96
 Increment in % 0.54 11.03 11.44 9.75
-ScanObjectNN OBJ + BG [SONN]
-DCP-v1 [DCP] 50.3 $\pm$5.25 48.2 $\pm$7.35 29.45 $\pm$5.34 32.15 $\pm$2.9
+ScanObjectNN OBJ + BG [41]
+DCP-v1 [42] 50.3 $\pm$5.25 48.2 $\pm$7.35 29.45 $\pm$5.34 32.15 $\pm$2.9
 MDCP-v1 (Ours) 62.2 $\pm$6.73 60.2 $\pm$7.54 37.25 $\pm$3.26 41.45 $\pm$3.21
 Increment in % 23.66 24.90 26.49 28.93
-DCP-v2 [DCP] 52.3 $\pm$6.92 55 $\pm$8.16 33.8 $\pm$4.31 36.25 $\pm$4.57
+DCP-v2 [42] 52.3 $\pm$6.92 55 $\pm$8.16 33.8 $\pm$4.31 36.25 $\pm$4.57
 MDCP-v2 (Ours) 58.8 $\pm$8.07 62.8 $\pm$6.22 41.9 $\pm$6.2 44.6 $\pm$4.1
 Increment in % 12.43 14.18 23.96 23.03
-ScanObjectNN PB75 [SONN]
-DCP-v1 [DCP] 46.1 $\pm$8.31 45.2 $\pm$9.37 25.2 $\pm$3.78 28.45 $\pm$4.5
+ScanObjectNN PB75 [41]
+DCP-v1 [42] 46.1 $\pm$8.31 45.2 $\pm$9.37 25.2 $\pm$3.78 28.45 $\pm$4.5
 MDCP-v1 (Ours) 54.8 $\pm$7.35 51.5 $\pm$7.08 34.05 $\pm$4.2 37.65 $\pm$4.69
 Increment in % 18.87 13.94 35.12 32.34
-DCP-v2 [DCP] 49.6 $\pm$6.64 49.6 $\pm$8.7 27.85 $\pm$3.75 32.35 $\pm$4.03
+DCP-v2 [42] 49.6 $\pm$6.64 49.6 $\pm$8.7 27.85 $\pm$3.75 32.35 $\pm$4.03
 MDCP-v2 (Ours) 54.7 $\pm$5.89 53.2 $\pm$9.46 34.45 $\pm$3.48 41.15 $\pm$4.87
 Increment in % 10.28 7.26 23.70 27.20
-Few-shot Classification.
-Table 2 evaluates the discriminative power of Mahalanobis distance.
-We compare features learned by DCP-v1/v2 [DCP] and MDCP-v1/v2 in few-shot classification tasks using the $k$-way, $m$-shot setting [CrossPoint].
-Models are trained for registration on ModelNet40 [modelnet] and then tested on few-shot classification for ModelNet40 and three ScanObjectNN [SONN] splits (OBJ ONLY, OBJ+BG, PB75).
-MDCP consistently outperforms DCP across all settings.
-Summarizing Table 2, MDCP-(v1,v2) achieves average accuracy improvement of $(22.98,6.52)$ on ModelNet40.
-For ScanObjectNN, improvements are: OBJ ONLY: $(12.68,8.19)$; OBJ+BG: $(25.99,18.40)$; PB75: $(25.06,17.11)$, [(a, b) denotes average accuracy increment (%) for MDCP-(v1,v2)].
-These results suggest two key points: 1) v1 (without transformer) outperforms v2, indicating transformers may struggle with limited data; 2) Even on unseen ScanObjectNN data, MDCP shows consistent improvement across splits.
+Few-shot Evaluation.
+The discriminative prowess of the proposed Mahalanobis distance is evaluated through Table 2, illustrating the performance of features learned by both DCP-(v1,v2) [42] and our proposed MDCP-(v1,v2) in the context of few-shot evaluation tasks.
+The evaluation adheres to a $k$-way, $m$-shot few-shot strategy detailed in  [1].
+The models are initially trained for the point-cloud registration task on the ModelNet40 dataset, and subsequently assessed for few-shot classification tasks on the ModelNet40 test set.
+Furthermore, evaluations extend to three distinct splits, namely OBJ ONLY, OBJ + BG, and PB75, from the ScanObjectNN [41] dataset.
+A striking observation is the substantial performance enhancement of the Mahalanobis version of DCP, which surpasses its original variant by a significant margin across all few-shot tasks.
+Summarizing the performance of Mahalanobis versions (v1,v2) as reported in Table 2, we increment across all few-shot tasks on the ModelNet40 [45] dataset of (22.98, 6.52).
+In the context of the ONLY OBJ split, an average increment of (12.68, 8.19) is recorded, while for the OBJ + BG split, the increments are (25.99, 18.40).
+Similarly, for the PB75 split, increments of (25.06, 17.11) are documented.
+Note: (a, b) denotes average accuracy increment in % for MDCP-(v1,v2).
+From the results, two insights emerge: Firstly, v1 without a transformer outperforms v2 with a transformer, suggesting that transformers may not generalize as effectively with limited data.
+Secondly, its worth noting that even when evaluated on unseen ScanObjectNN [41] data, which was not included during the training phase, a consistent increment is observed across all splits.
 5.3 Endurance Test
-Table 3: Quantitative comparison of MDCP-(v1,v2) robustness against original on ModelNet40 [modelnet] for low-density point cloud registration.
-We evaluate across various settings.
-MDCP consistently outperforms its original counterpart (bold denotes best).
-Lower is better for RMSE(t) (reported in $10^{-2}$ units).
-Note: -Geo refers to the vectorized Floyd-Warshall geodesic-graph version.
-Target Low Density ModelNet40 [modelnet]
+Table 3: Quantitative comparison to evaluate the robustness of the proposed MDCP-(v1,v2) against its original variants, benchmarked on the ModelNet40 [45] dataset.
+This comparison is conducted across various settings for the point-cloud registration task, particularly when the target and source point cloud exhibit low point density.
+Throughout the comparison, our approach consistently demonstrates evident superiority in performance.
+The bold typeface denotes the best results.
+It is important to note that all evaluation metrics follow a “lower is better” criterion, and the RMSE(t) metric is reported in units of $10^{-2}$.
+Note: -Geo refers to Flyod-Warshall geodesic-graph version.
+Target Low Density ModelNet40 [45]
 Unseen Data Unseen Category Noise
 RMSE(R) RMSE(t) RMSE(R) RMSE(t) RMSE(R) RMSE(t)
-1024 DCP-v1 [DCP] 42.0026 10.865 8.3181 1.2206 32.0181 0.9616
+1024 DCP-v1 [42] 42.0026 10.865 8.3181 1.2206 32.0181 0.9616
 DCP-v1-Geo 38.1172 10.863 6.0088 1.2806 39.9996 0.9801
 MDCP-v1 32.6880 10.977 5.8425 1.2096 24.9222 1.0062
-DCP-v2 [DCP] 68.0209 3.6777 14.9977 2.7765 66.9599 3.0532
+DCP-v2 [42] 68.0209 3.6777 14.9977 2.7765 66.9599 3.0532
 DCP-v2-Geo 44.6676 3.7786 14.7762 8.1678 48.2278 2.9987
 MDCP-v2 34.9602 3.3871 14.3836 6.1023 25.8595 2.8308
-512 DCP-v1 [DCP] 54.5579 1.6058 8.31497 1.6812 50.4602 1.4233
+512 DCP-v1 [42] 54.5579 1.6058 8.31497 1.6812 50.4602 1.4233
 DCP-v1-Geo 48.1628 1.7898 8.1065 1.6711 49.7866 1.5661
 MDCP-v1 42.3370 1.4918 8.4701 1.6028 38.0739 1.3168
-DCP-v2 [DCP] 80.8099 8.9802 23.8370 3.7857 77.4393 8.7660
+DCP-v2 [42] 80.8099 8.9802 23.8370 3.7857 77.4393 8.7660
 DCP-v2-Geo 77.8756 6.7752 21.8861 3.6544 48.2165 8.7765
 MDCP-v2 47.3271 3.9172 28.7305 8.4551 44.3240 3.4637
-Source Low Density ModelNet40 [modelnet]
-1024 DCP-v1 [DCP] 43.2637 1.1146 8.304669 1.228 38.6104 1.0306
+Source Low Density ModelNet40 [45]
+1024 DCP-v1 [42] 43.2637 1.1146 8.304669 1.228 38.6104 1.0306
 DCP-v1-Geo 36.6378 1.1123 8.0011 1.2761 39.1654 1.1756
 MDCP-v1 31.3686 1.1035 8.3409 1.2247 26.9862 1.0158
-DCP-v2 [DCP] 65.4777 3.1448 16.1993 1.9267 61.1659 2.6327
+DCP-v2 [42] 65.4777 3.1448 16.1993 1.9267 61.1659 2.6327
 DCP-v2-Geo 57.1998 3.0098 14.7176 6.7861 49.1109 2.9987
 MDCP-v2 38.6540 2.8759 7.4574 5.6750 27.1185 2.6114
-512 DCP-v1 [DCP] 55.2426 1.5553 8.4311 1.4074 51.7135 1.3908
+512 DCP-v1 [42] 55.2426 1.5553 8.4311 1.4074 51.7135 1.3908
 DCP-v1-Geo 38.1172 1.0863 6.0088 1.2806 39.9996 0.9801
 MDCP-v1 42.4809 1.4700 7.8250 1.1089 39.6634 1.3272
-DCP-v2 [DCP] 71.8377 7.2431 22.3028 4.8566 71.1976 6.9627
+DCP-v2 [42] 71.8377 7.2431 22.3028 4.8566 71.1976 6.9627
 DCP-v2-Geo 44.6676 3.7698 26.8861 4.9908 65.9981 5.0098
@@
 Efficiency towards Point-density.
-Table 3 evaluates MDCP’s robustness against varying point densities, simulating real-world sensor data [survey].
+Table 3 evaluates MDCP’s robustness against varying point densities, simulating real-world sensor data [23].
 One point cloud (source or target) is downsampled by half (e.g., 2048 → 1024).
 Evaluation metrics use the same number of points for both clouds (e.g., source with 1024 points and target with 2048 points are used to compute transformation, then applied to original points for error calculation).
-Figure 3 compares DCP-v1 [DCP] and MDCP-v1.
+Figure 3 compares DCP-v1 [42] and MDCP-v1.
 When densities are similar (man’s face), both perform well.
@@
 Note: Source point cloud is a transformed version of $g(Y)$, and the goal is to estimate $g$ for perfect alignment with target $X$.
-Table 4: Quantitative comparison of proposed MDeepUME’s robustness on ModelNet40 [modelnet], FAUST [FAUST], and Stanford 3D datasets.
-We evaluate across various noise, benchmarked by authors in DeepUME [DeepUME].
-MDeepUME consistently outperforms DeepUME (bold denotes best results).
-Lower is better for all metrics: RMSE(t), chamfer distance (CD) [ChamferDistance], and Hausdorff distance (HD) [HausdorffDistance] (reported in $10^{-2}$ units).
-Noise-types DeepUME [DeepUME] MDeepUME
+Efficiency towards Point-density.
+Table 3 illustrates the robustness of the proposed Mahalanobis version of DCP [42] in the context of point-cloud registration tasks that simulate real-world scenarios.
+These scenarios involve two point clouds originating from distinct sensors [23], potentially exhibiting differing point densities.
+In this experimental setup, one of the two point clouds, either the source or target, is downsampled by half relative to the other.
+As depicted in Table 3, the value 1024 signifies that either the target or source point cloud is downsampled to half its original size (e.g., 2048 $\to$ 1024).
+Note: To compute evaluation metrics, both source and target point clouds are ensured to have the same number of points.
+Thus, rigid-transformation parameters are computed using the downsampled point cloud and applied to the original-scale points (e.g., $X_{2048}$ and $Y_{1024}$ are used to generate transformation parameters $g$, which are then utilized to evaluate RMSE($X_{2048}$, $g(Y_{2048})$)).
+This scenario is effectively illustrated in Figure 3, where the performance comparison between DCP-v1 [42] and MDCP-v1 is particularly insightful.
+In cases where the point clouds have same density, such as in the highlighted regions depicting a man’s face, DCP-v1 and MDCP-v1 exhibit relatively close performance.
+However, the subtleties become more pronounced when subjected to the aforementioned settings.
+Specifically, DCP-v1 struggles to accurately estimate rigid transformations, leading to subtle angular discrepancies as evident from the highlighted regions.
+In contrast, the proposed MDCP-v1 showcases a robustness to varying point densities, as demonstrated through the highlighted regions portraying a chair.
+Note: The source point cloud is a transformed version of $g(Y)$, wherein the objective is to estimate the rigid transformation parameters $g$ that align it perfectly with the target point cloud $X$.
+Table 4: Quantitative comparison to evaluate the robustness of the proposed MDeepUME against its original variant, benchmarked on the ModelNet40 [45], FAUST [7] and Stanford 3D repository dataset.
+This comparison is conducted across various noise types proposed in DeepUME [27] for the point-cloud registration task.
+Throughout the comparison, our approach consistently demonstrates evident superiority in performance.
+The bold typeface is utilized to denote the best results.
+It is important to note that all evaluation metrics follow a “lower is better” criterion, and the RMSE(t), chamfer-distance [3] (CD), and hausdorff-distance [14] (HD) metric are reported in units of $10^{-2}$.
+Noise-types DeepUME [27] MDeepUME
 CD HD RMSE(R) RMSE(t) CD HD RMSE(R) RMSE(t)
@@
 Efficiency towards Various Noise Types.
-Table 4 evaluates the robustness of MDeepUME against various noise types from DeepUME [DeepUME] (Bernoulli, Gaussian, Sampling, Zero-Intersection).
-We benchmark on ModelNet40 [modelnet] (training and evaluation), FAUST [FAUST], and Stanford3D (testing only).
-Training settings follow the original paper [DeepUME] ( [^3] https://github.com/langnatalie/DeepUME).
-MDeepUME outperforms DeepUME across datasets and noise types in terms of Chamfer distance (CD) [ChamferDistance], Hausdorff distance (HD) [HausdorffDistance], and root mean squared error in rotation (RMSE(R)).
-However, limitations are observed in translation accuracy (RMSE(t)) due to DeepUME’s inherent projection into SO(3)-invariant space, where Mahalanobis distance is less effective.
+Table 4 illustrates the resilience of the proposed Mahalanobis version of DeepUME [27] in the context of point cloud registration tasks subjected to various noise sampling strategies, as outlined in DeepUME [27].
+Our evaluation encompasses diverse noise types, including Bernoulli, Gaussian, Sampling, and Zero-Intersection, each meticulously detailed within the framework of DeepUME [27].
+The benchmarking is conducted on the ModelNet40 [45], FAUST [7], and Stanford-3D Repository datasets.
+While the latter two datasets are exclusively utilized for testing, ModelNet40 is employed for comprehensive evaluation.
+All training-testing settings [^3] https://github.com/langnatalie/DeepUME are considered wrt original paper [27].
+The outcomes presented in Table 4 substantiate the superior performance of the proposed MDeepUME over its original counterpart across various datasets and noise types.
+This observation holds true for a multitude of evaluation metrics, notably including Chamfer distance [3] (CD), Hausdorff distance [14] (HD), and root mean squared error in rotation (RMSE(R)).
+However, it is worth noting that MDeepUME demonstrates certain limitations, particularly in estimating translation accuracy, as evident from the RMSE(t) values in Table 4.
+This outcome is attributed to DeepUME’s intrinsic projection of points into an $\mathcal{SO}(3)$-invariant space, wherein the potential of Mahalanobis distance cannot be fully harnessed.
+5.4 Limitations
+While our approach harnesses the prowess of Mahalanobis distance for enhanced feature matching and point cloud registration, it is important to acknowledge its limitations.
+Mahalanobis distance computation critically depends on the accurate estimation of the covariance matrix.
+When the covariance matrix is ill-conditioned, the precision may become compromised.
+Although we mitigate this concern by introducing a small bias term ($10^{-5}$) to the diagonal of the covariance matrix.
+This limitation is evident in noise benchmark evaluation for DeepUME, where the Mahalanobis version fails in accurately estimating translation parameters.
+Despite these limitations, our approach significantly contributes to address key challenges in point cloud registration.
 6 Conclusion
-Our paper introduces Mahalanobis k-NN, a statistical framework for improving point cloud registration in learning-based methods.
+Our paper introduced Mahalanobis k-NN, a statistical framework for improving point cloud registration in learning-based methods.
 It addresses challenges like feature matching due to variations in point cloud densities.
@@
 It integrates seamlessly into local graph-based point cloud analysis methods like DCP and DeepUME, achieving state-of-the-art performance on benchmark datasets.
-We also demonstrate that registered point cloud features possess discriminative capabilities, leading to significant improvements in few-shot classification tasks.
-Mahalanobis k-NN offers exceptional surface awareness and outlier robustness, benefiting tasks like segmentation, decomposition, and normal estimation.
-Additionally, it serves as a surficial geometry-based error metric, enhancing tasks like generation and upsampling.
-Its versatility makes it a crucial component in advancing point cloud processing, warranting further benchmarking to highlight its significance.
-[…]
+We also demonstrated that registered point cloud features possess discriminative capabilities, leading to significant improvements in few-shot classification tasks.
+Mahalanobis k-NN offers surface awareness and outlier robustness.
+Overall, our proposed method offers a comprehensive and effective solution for point cloud registration, providing superior results, and versatility across various benchmarking test-beds.